# VENI, VINDy, VICI – a variational reduced-order modeling framework with uncertainty quantification

### v1

@@ 3.1 Numerical example I: Rössler System
In order to assess the potential of our novel VINDy approach alone, we first test the methodology on a low-dimensional system. Consequently, no autoencoder is present and the VINDy method directly operates on the original system description. We consider the Rössler system for this validation. It is a set of ordinary differential equations defined as
@@
where $\bm{z}=[z_{1},z_{2},z_{3}]^{\intercal}$ is the vector of the system states and $\bm{\alpha}=[\alpha_{1},\alpha_{2},\alpha_{3}]^{\intercal}$ contains the system parameters.
Dataset
The data used to train our model consists of multiple simulations with parameters $\alpha_{1}=0.2,$ $\alpha_{2}=0.2,$ and $\alpha_{3}=5.7$ as well as normal randomly distributed initial conditions $\bm{z}(t_{0})\sim\mathcal{N}([-5,-5,0]^{\intercal},\sigma^{2}\bm{I})$ with a standard deviation of $\sigma=2.25$. We consider two sources of uncertainty: measurement noise and model noise. For the former we apply multiplicative noise to the state measurements, ı.e. $\bm{z}_{\text{noise}}(t)=\varepsilon\bm{z}(t)$ with $\varepsilon\sim\mathcal{L}\mathcal{N}(\bm{0},\lambda_{\varepsilon}*\bm{I})$ drawn from a log-normal distribution individually for every sample. The time derivatives are numerically computed from those noisy samples. The model uncertainty is realized as noise in the model parameters which are drawn from a normal distribution $\bm{\alpha}\sim\mathcal{N}([\alpha_{1},\alpha_{2},\alpha_{3}]^{\intercal},[\sigma_{\alpha_{1}}^{2},\sigma_{\alpha_{2}}^{2},\sigma_{\alpha_{3}}^{2}]^{\intercal})$ centered around the correct values with standard deviations $\sigma_{\alpha_{1}}=\lambda_{\alpha}{\alpha_{1}}$$\sigma_{\alpha_{2}}=\lambda_{\alpha}{\alpha_{2}}$, and $\lambda_{\alpha}\sigma_{\alpha_{3}}={\alpha_{3}}$. The test data consists of a set of simulation that differs in its initial conditions from the training data and uses the correct parameter values.
@@
For the VINDy layer, we use a polynomial library $\bm{\Theta}(\bm{z})\in\mathbb{R}^{10}$ of order 2 including bias and interactions with Laplacian priors $\bm{\Xi}_{ij}\sim\mathcal{L}(0,1)$ for the corresponding coefficients. We train the model over 2000 epochs using early stopping. At the end of the training all coefficients $\bm{\Xi}_{ij}$ with a probability density function above a threshold of $pdf_{\bm{\Xi}_{ij}}(0)>5$ are set to zero.
Experiments and Results
We perform six experiments to measure the performance of VINDy: (a) One on noise-free data, two with model noise ((b) moderate $\lambda_{\alpha}=0.1$, (c) high $\lambda_{\alpha}=0.2$), two with measurement noise ((d) moderate $\lambda_{\varepsilon}=0.05$, (e) high $\lambda_{\varepsilon}=0.1$), and (f) one where both sources of uncertainty are taken into account. The impact of the respective noise on the training data can be seen in Fig. 4. For each training data set, we apply VINDy and sample from the resulting coefficient distributions to perform a sample-based UQ on the time evolution of the system states. As shown in Fig. 4, the resulting trajectories for no noise (a) and moderate model noise (b) follow the ground-truth quite well and capture the system dynamics. As expected, higher model noise leads to broader uncertainty bounds around the predicted mean trajectories. Nevertheless, even for high model noise (c) the identified system is able to follow the reference. This also holds true for moderate measurement noise (d) and the approach is only pushed to its limits for high measurement noise (e) and the combination of both noises (f). In those cases the state trajectories differ from the ground truth. However, even in these cases, the method extracts essential knowledge from the data, which becomes apparent upon looking at the identified coefficient distributions.
Figure 4: Temporal evolution of the Rössler system states including UQ for a test scenario on various noise scenarios. The corresponding training data for each scenario is shown at the right of the state trajectories.
As shown in Fig. 5, our method keeps all the important terms that originally appear in the underlying equation without relevance what kind of uncertainty is present in the data, i.e. for all noise levels. At the same time, most of the other irrelevant coefficients are centered around zero. For moderate noise level, the resulting coefficient distributions center around the reference value with high confidence, whereas higher noise levels lead to more conservative estimates and wider distributions.
Figure 5: Distributions of coefficients identified for the Rössler system based on noise-free and noisy training data. The correct values are marked with a triangle at the bottom of each axis. All values that occur in the original equations are highlighted with red axes.
3.2 Numerical example II: Reaction-diffusion problem
We consider a reaction-diffusion system governed by equations
@@ 3.2 Numerical example II: Reaction-diffusion problem
where $\mu=1.0$ and $d_{1}=d_{2}=0.01$ are coefficients that respectively regulate the reaction and diffusion behaviors of the system. The problem is defined over a spatial domain $[-L,L]^{2}$ for $L=10$ and a time span $t\in[0,T]$ for $T=40$, periodic boundary conditions are prescribed, and the initial condition is defined as
$u(x,y,0;\beta)=v(x,y,0;\beta)=\tanh{\left(\beta\sqrt{x^{2}+y^{2}}\cos{\left((x+iy)-\beta\sqrt{x^{2}+y^{2}}\right)}\right)}\,,$ (20)
depending on the parameter $\beta\in\mathcal{P}=[0.7,1.1]$. System (19) generates spiral waves, which represent an attracting limit cycle in the state space [41] and whose radius is modulated by the parameter $\beta$.
Our goal is to create a generative model to compute the entire space–time solution for a new instance of parameter $\beta$.
Dataset
A grid of $N_{\beta}=20$ values for the parameter $\beta$ equispaced over $\mathcal{P}$ is considered. The parameter values are randomly partitioned in training and testing subsets, $\mathcal{P}_{\text{train}}$ and $\mathcal{P}_{\text{test}}$, such that $N_{\beta}^{\text{train}}=|\mathcal{P}_{\text{train}}|=16$ and $N_{\beta}^{\text{test}}=|\mathcal{P}_{\text{test}}|=4$. For each choice of $\beta$, numerical solutions for $u$ and $v$ are computed by solving the PDEs (19) with initial condition (20) by using the Fourier spectral method [42] with time step $\Delta t=0.05$ on an equispaced spatial grid with spatial step $\Delta h=0.4$. For the training set, solutions are computed over a limited time window $T_{\text{train}}=20$, while testing data reach final time $T=40$. Moreover, training data are corrupted by $20\%$ log-normal multiplicative noise to simulate measurement noise. Training and testing snapshots are stacked respectively in matrices $\mathbf{X}_{\text{train}}\in\mathbb{R}^{N_{t}^{\text{train}}N_{\beta}^{\text{train}}\times N}$ and $\mathbf{X}_{\text{test}}\in\mathbb{R}^{N_{t}^{\text{test}}N_{\beta}^{\text{test}}\times N}$, where $N_{t}^{\text{train}}=400$ and $N_{t}^{\text{test}}=800$ are the number of samples for training and testing sets respectively, and $N=2500$ is the number of spatial degrees of freedom. We preliminary reduce system dimension to $N_{\text{POD}}=32$ by projecting both training and testing data onto the reduced basis obtained using POD on $\mathbf{X}_{\text{train}}$. Projected data are denoted as $\tilde{\mathbf{X}}_{\text{train}}\in\mathbb{R}^{N_{t}^{\text{train}}N_{\beta}^{\text{train}}\times N_{\text{POD}}}$ and $\tilde{\mathbf{X}}_{\text{test}}\in\mathbb{R}^{N_{t}^{\text{test}}N_{\beta}^{\text{test}}\times N_{\text{POD}}}$.
@@
Regarding the VENI structure, the encoder consists out of a feed-forward neural network with 3 hidden dense layer of 32, 16 and 8 units, while the decoder network has a symmetrical structure. Spiral waves generated by the system (19) can be approximately captured by two oscillating spatial modes [17], thus we set the number of latent variables $n=2$. For VINDy, we employ a set of polynomial functions of the latent variables $\mathbf{z}=\left[z_{1},z_{2}\right]^{\top}$ up to the third degree with interaction and bias. No parametric dependency is included in the library, as parameter $\beta$ affects solely the initial conditions. Regarding the choice of prior distributions on the latent variables and on the VINDy coefficients, we consider gaussian and laplacian distributions, respectively, namely $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\bm{I})$ and $\bm{\Xi}_{ij}\sim\mathcal{L}(0,1)$ for each entry.
Experiments and Results
@@
The offline training is performed on the dataset $\{\tilde{\mathbf{X}}_{\text{train}},\dot{\tilde{\mathbf{X}}}_{\text{train}}\}$. The resulting posterior distributions for the model coefficients are depicted in Fig. 6. The primary dynamics of this phenomenon exhibit the characteristics of a linear oscillator, which our method successfully captures by identifying the linear coefficients as significant terms with high confidence. These coefficients are distinguished by a posterior mean significantly deviating from zero and a notably narrow variability, confirming their critical role in the observed dynamics. Moreover, our approach encourages parsimony in the dynamics by effectively setting to zero mean most nonlinear terms. Fig. 6 also presents the phase space representation in the identified latent coordinates, showcasing the oscillatory pattern and the attracting limit cycle. This demonstrates not only the effectiveness of our method but also its ability to preserve interpretable insights into the latent space.
@@
To assess the accuracy and reliability of our method in making predictions, we evaluate its performance on the test set using the VICI approach. For initial conditions not seen during training and corresponding to test parameter instances, we predict the approximate solutions up to a final time of $T=40$. These predictions are compared with the noise-free, numerical values ${\mathbf{X_{\text{test}}}}$. Fig. 7 illustrates the mean and standard deviation over $m=100$ realizations of the predicted reconstruction of the entire solution field. We note how the method is able to accurately detect the spatial dynamics, even at extrapolating time instances far from the training range. To highlight the temporal extrapolation capability of our method and to illustrate how uncertainty propagates during the reconstruction phase, we also present in Fig. 8 the temporal evolution of the predicted variables at the different reconstruction levels.
Our method demonstrates great predictive performance, accurately approximating the system’s evolution in a time window twice longer than the training range.
@@
Figure 6: Results of the offline training, consisting of the VENI and VINDy steps, for the reaction-diffusion problem. In the VINDy box, we display the identified posterior distributions of the coefficients for the latent dynamic system. The rows represent the candidate terms considered for the dynamics, and the columns indicate which latent dynamic equation they belong to. We note that the mixed linear terms, which are the dominant terms for the observed dynamics, are accurately identified as significant. Most nonlinear terms contribute a mean of zero, and we observe that the uncertainty of the coefficients increases with the degree of the polynomial. In the VENI box, we highlight how the method is capable of learning an encoding function from noisy data, resulting in two latent variables that exhibit the expected oscillatory behavior.
Figure 7: Comparison of solution fields in the reaction-diffusion example between the approximation by the proposed method and the numerical reference solution truth. The snapshots refer to one training time instances $t=15$ and two extrapolated time instances $t\in\{25,35\}$ (being $T_{\text{train}}=20$) for two testing parameter $\beta=\{0.93,0.84\}$. We report the prdictions of the mean and standard deviation of the solutions predicted by the VICI procedure.
Figure 8: Forward uncertainty quantification at the different reduction levels, performed with the VICI procedure. The dashed line at $T_{\text{train}}=20$ indicates the end time of training data coverage. The red line indicates the noise-free reference solution, the dashed blue line the predicted mean, and the shaded region indicates the uncertainty bounds.
3.3 Numerical example III: Beam MEMS resonator
We consider a straight beam MEMS resonator, excited at resonance, which could be represented as double clamped beam. The illustration and the properties of the device are reported in Fig. 9.
Figure 9: Schematic representation of the beam MEMS resonator with the mesh used in the FOM simulations. The beam has length $L=~{}1000$$\mu$m with a rectangular cross-section of dimensions 10 $\mu$m $\times$24 $\mu$m , made of isotropic polysilicon [43], with density $\rho=~{}2330$ Kg/m3, Young modulus $E=167$ GPa and Poisson coefficient $\nu=0.22$. The first bending eigenfrequency is $\omega_{0}=0.5475\,\text{rad}/\mu\text{s}$. Dirichlet boundaries are highlighted.
The governing PDEs for this solid mechanics problem in large transformations [44] are formulated in terms of the displacement field $\mathbf{u}$ as follows:
@@ 3.3 Numerical example III: Beam MEMS resonator
where $\Omega_{0}$ is the domain occupied by the device in the undeformed configuration, described by material coordinates $\varkappa$, and $T=1755\,\mu\text{s}$.Eq. (21a) expresses the conservation of momentum where $\rho_{0}$ is the initial density, $\mathbf{P}$ is the first Piola-Kirchhoff stress and $\mathbf{B}$ are the external body forces. In this application we consider a periodic body load proportional to the first vibrational eigenfunction, with amplitude and frequency given by the parameters $F$ and $\omega$, respectively. The input parameter vector hence becomes $\bm{\beta}=[\omega,F]$, defined over a closed and bounded set of dimension 2. Eq. (21b) and (21c) define the homogeneous boundary conditions on the Neumann and Dirichlet and boundaries denoted as $\partial\Omega_{N}$ and $\partial\Omega_{D}$, respectively, and where $\mathbf{N}$ denotes the outward-directed unit normal on the boundary of $\Omega$. For more details regarding the application, we refer to [19].
Dataset
We consider $N_{\beta}=56$ parameter instances, corresponding to 28 values for the frequency $\omega$, selected in the range $[0.526,0.564]\,\text{rad}/\mu\text{s}$ with a finer sampling around the natural frequency $\omega_{0}$, and for two values of the forcing amplitude $F\in\{0.125,0.250\}\,\mu\text{N}$. In total, $N_{\beta}^{\text{train}}=28$ parameter instances are randomly selected to construct the training set and the remaining $N_{\beta}^{\text{test}}=28$ are retained for testing. Numerical solutions for the displacement are computed for each parameter configuration, by discretizing the system (21) using the finite element method over a spatial mesh consisting of $N_{\text{dof}}=7821$ degrees of freedom and using finite element method as in [19] of which only 5964 are nonzero. As we are modeling the beam as a second-order system, we require to have access or compute the second time derivatives $\ddot{\bm{x}}$ for training in addition to the system states $\bm{x}$ and their time derivatives $\dot{\bm{x}}$.
@@
[…]
As in the previous example, the data (including displacements, velocities, and accelerations) are preliminary projected onto a POD basis of $N=N_{\text{POD}}=3$ modes (extracted from the noisy displacement training data). Finally, snapshots of the displacement are collected in $\tilde{\mathbf{X}}_{\text{train}}\in\mathbb{R}^{N_{t}^{\text{train}}N_{\beta}^{\text{train}}\times N}$ and $\tilde{\mathbf{X}}_{\text{test}}\in\mathbb{R}^{N_{t}^{\text{train}}N_{\beta}^{\text{train}}\times N}$ matrices, where $N_{t}^{\text{train}}=14001$ and $N_{t}^{\text{test}}=22499$ are the corresponding numbers of time instances. Analogously, velocities are collected in $\dot{\tilde{\mathbf{X}}}_{\text{train}}$ and accelerations in $\ddot{\tilde{\mathbf{X}}}_{\text{train}}$.
@@
The encoder is composed of three hidden layers using 32 neurons each. The decoder has symmetrical structure. Although the dynamics of this system lies on a two dimensional manifold in the phase space, we manage to reduce the dimensionality to a single latent variable $n=1$, as the dependence on the velocity is the minimal for the master bending mode that we inspect herein. As in (21), structural mechanics problems are often governed by second-order governing equations. Therefore we model the reduced dynamics with a second-order ODE, by augmenting the reduce states with the first derivatives $\dot{z}$:
$\frac{\text{d}}{\text{dt}}\begin{bmatrix}\;z\;\\ \;\dot{z}\;\end{bmatrix}=\mathbf{f}\left(z,\dot{z};\bm{\beta}\right)\approx\begin{bmatrix}\;\dot{z}\;\\ \;\bm{\Theta}(z,\dot{z};\bm{\beta})\bm{\Xi}\;\end{bmatrix}.$ (22)
The corresponding accelerations at latent level $\ddot{\bm{z}}$ can be derived using chain rule as it is done for the velocities $\dot{\bm{z}}$.
Note that, in case we did not want to exploit any prior knowledge on the second-order structure of the problem, we could have equivalently modelled the problem by considering $n=2$ latent variables and identifying a first-order ODE system as in the previous example. The library $\bm{\Theta}$ features polynomials up to the third degree with respect to both $\bm{z}$ and $\dot{\bm{z}}$. Moreover, since the beam resonator is forced harmonically with amplitude and frequency parameterized by $F$ and $\omega$, we incorporate the parameter dependency by including the library term $F\cos(\omega t)$. We set gaussian priors on the reduced states $[\bm{z},\dot{\bm{z}}]^{\top}\sim\mathcal{N}(\mathbf{0},\bm{I})$ and laplacian prior on each of the VINDy coefficient distributions $\bm{\Xi}_{ij}\sim\mathcal{L}(0,1)$. Training is performed for 2500 epochs, and the weights of the networks that performed best in terms of total loss on a validation subset of the training data used for inference. After training we apply PDF zero tresholding, see Section 2.5.1, to only keep the terms that are important to describe the dynamics. In particular, all terms with a $pdf_{\bm{\Xi}_{ij}}(0)>1.9$ are neglected in the model.
3.3.1 Results

### v2

@@ 3.1 Rössler System: Identification of a low-dimensional chaotic systems under varying noise-levels
In order to assess the potential of our novel VINDy approach alone (without autoencoder), we present results for a low-dimensional chaotic system. Therefore, we consider the Rössler system
@@
where $\bm{z}=[z_{1},z_{2},z_{3}]^{\intercal}$ is the vector of the system states and $\bm{\alpha}=[\alpha_{1}=0.2,\alpha_{2}=0.2,\alpha_{3}=5.7]^{\intercal}$ contains the system parameters. Model hyperparameters and dataset configurations are summarized in the A.4 in Table 1-2.
Data generation
The data used to train our model consists of 30 simulations of the Rössler System (17) with normal randomly distributed initial conditions $\bm{z}(t_{0})\sim\mathcal{N}([-5,-5,0]^{\intercal},\sigma^{2}\bm{I})$ with $\sigma=2.25$. Each simulation has a time horizon of $T=24$ and consists out of $N_{t}^{\text{train}}=2000$ time steps. We consider two sources of uncertainty: measurement noise and model noise. For the former we apply multiplicative noise to the state measurements, ı.e. $\bm{z}_{\text{noise}}(t)=\bm{\varepsilon}*\bm{z}(t)$ with $\bm{\varepsilon}\sim\mathcal{L}\mathcal{N}(\bm{0},\lambda_{\bm{\varepsilon}}\bm{I})$ drawn from a log-normal distribution individually for every sample. The time derivatives are numerically computed from those noisy samples. The model uncertainty is realized as noise in the model parameters which are drawn from a normal distribution $\bm{\alpha}\sim\mathcal{N}\left([\alpha_{1},\alpha_{2},\alpha_{3}]^{\intercal},\text{diag}([\sigma_{\alpha_{1}}^{2},\sigma_{\alpha_{2}}^{2},\sigma_{\alpha_{3}}^{2}]^{\intercal})\right)$ centered around the correct values with standard deviations $\sigma_{\alpha_{1}}=\lambda_{\alpha}{\alpha_{1}}$$\sigma_{\alpha_{2}}=\lambda_{\alpha}{\alpha_{2}}$, and $\sigma_{\alpha_{3}}=\lambda_{\alpha}{\alpha_{3}}$. For the experiments in Fig. 3, we applied $\lambda_{\alpha}=0.1$ for model noise and $\lambda_{\bm{\varepsilon}}=0.05$ respectively, while $\lambda_{\alpha}=0.2$ and $\lambda_{\bm{\varepsilon}}=0.1$ are used for the combined noise scenario. The test data consist of simulations that differ from the training data in their initial conditions, while using the correct parameter values. We consider two testing scenarios: initial conditions (ICs) sampled within the training regime, and out-of-distribution (OOD) ICs, drawn from a Gaussian distribution $\bm{z}(t_{0})\sim\mathcal{N}\!\left([-10,\,-10,\,20]^{\intercal},\,\sigma^{2}\bm{I}\right).$
@@
For the VINDy layer, we use a polynomial library of second degree, including bias and interactions, for a total of $r=10$ candidate features. We consider Laplacian priors $\bm{\Xi}_{ij}\sim\mathcal{L}(0,1)$ for the corresponding coefficients. The model is trained for 500 epochs. At the end of the training, all coefficients $\bm{\Xi}_{ij}$ with a probability density function above a threshold of $pdf_{\bm{\Xi}_{ij}}(0)>5$ are removed.
Results
We identify governing equations using VINDy with a polynomial library of candidate functions under model and measurement noise. The impact of the respective noise on the training data can be seen in Fig. 3. The resulting trajectories under model noise closely match the ground truth, and a similar agreement is observed for measurement noise. Even in the OOD case, the predicted trajectories remain well aligned, albeit with broader uncertainty intervals. The approach only begins to reach its limits when high model and measurement noise are present simultaneously. However, even in the latter case, the method identifies all the essential terms that originally appear in the governing equations, regardless of the type of uncertainty hampering the data. At the same time, the irrelevant coefficients are centered around zero with corresponding high values of their PDF. For moderate noise levels, the resulting coefficient distributions center around the reference value with high confidence, whereas higher noise levels lead to more conservative estimates and wider distributions.
Figure 3: Approximation results for the Rössler system. Different types of noisy data (each represented by a different color) a) are generated by the Rössler system, b) serve as basis for identifying the distributions of coefficients, and c) are compared against the true values (indicated by triangles at the bottom of each axis). Coefficients appearing in the original equations are emphasized with bold red axes, while the PDF threshold is marked by a dashed line. The resulting temporal evolutions of the identified systems, including UQ, are shown in d) for an initial condition (IC) within the training regime and in e) for an IC outside the training domain (out-of-distribution inference).
3.2 Identification of a reaction-diffusion problem
We consider a reaction-diffusion system governed by
@@ 3.2 Identification of a reaction-diffusion problem
where $\mu=1.0$ and $d_{1}=d_{2}=0.01$ are coefficients that respectively regulate the reaction and diffusion behaviors of the system. The problem is defined over a spatial domain $L,L]^{2}$ for $L=10$ and a time span $t\in[0,T]$ for $T=40$, periodic boundary conditions are prescribed, and the initial condition is defined as
[…]
depending on the parameter $\beta\in[0.7,1.1]$. System (18) generates spiral waves, which represent an attracting limit cycle in the state space [39] whose radius is modulated by the parameter $\beta$. Our goal is to create a generative model to compute the entire space–time solution for a new instance of the parameter $\beta$. A.4 summarizes the model hyperparameters and data configuration parameters in Table 1-2.
Data generation
A grid of $N_{\beta}=20$ values is considered for the parameter $\beta$ equispaced with $[0.7,1.1]$. The parameter values are randomly partitioned into training and testing subsets such that $N_{\beta}^{\text{train}}=16$ and $N_{\beta}^{\text{test}}=4$. For each choice of $\beta$, numerical solutions for $u$ and $v$ are calculated by solving the PDE system (18) with initial condition (19) by using the Fourier spectral method [40] with time step $\Delta t=0.05$ on an equispaced spatial grid with spatial step $\Delta h=0.4$. For the training set, the solutions are calculated for a limited time window to $T_{\text{train}}=20$, while the testing data reach the final time $T=40$. In addition, training data are corrupted by $20\%$ log-normal multiplicative noise to simulate measurement noise. Training and testing snapshots are stacked in matrices $\mathbf{X}_{\text{train}}\in\mathbb{R}^{N_{t}^{\text{train}}N_{\beta}^{\text{train}}\times N}$ and $\mathbf{X}_{\text{test}}\in\mathbb{R}^{N_{t}^{\text{test}}N_{\beta}^{\text{test}}\times N}$, where $N_{t}^{\text{train}}=400$ and $N_{t}^{\text{test}}=800$ are the number of samples for training and testing sets respectively, and $N=2500$ is the number of spatial degrees of freedom. We preliminary reduce system dimension to $N_{\text{POD}}=32$ by projecting both training and testing data onto the reduced basis obtained using POD on $\mathbf{X}_{\text{train}}$.
@@
Regarding the VENI structure, the encoder consists out of a feed-forward neural network with 3 hidden dense layers of 32, 16 and 8 units, while the decoder network has a symmetrical structure. Spiral waves generated by the system (18) can be approximately captured by two oscillating spatial modes [10], thus setting the number of latent variables $n=2$. For VINDy, we employ a library of polynomial functions of the latent variables $\mathbf{z}=\left[z_{1},z_{2}\right]^{\top}$ up to the third degree with interaction and bias. No parametric dependency is included in the library, as parameter $\beta$ affects only the initial conditions. Regarding the choice of prior distributions on the latent variables and on the VINDy coefficients, we consider Gaussian and Laplacian distributions, respectively, namely $\bm{z}\sim\mathcal{N}(\mathbf{0},\bm{I})$ and $\Xi_{ij}\sim\mathcal{L}(0,1)$ for each entry.
Results
@@
The posterior distributions identified by our method for the model coefficients are depicted in Fig. 4. We note that the mixed linear terms, which are the dominant terms for the observed oscillatory dynamics, are accurately identified as significant with high confidence. In addition, our approach promotes sparsity in the dynamics, since most nonlinear terms are represented by distributions with high probability density function values at zero. Fig. 4 also presents the phase space representation in the identified latent coordinates, showcasing the oscillatory pattern and the attracting limit cycle. This demonstrates our method’s ability to preserve interpretable insights into the latent space.
@@
To assess the accuracy and reliability of VICI, we evaluate its performance for initial conditions and test parameter instances unseen during training. We predict the approximate solutions up to a final time of $T=40$, while we only use training data up to $T_{\text{train}}=20$. In Fig. 4, we illustrate the mean and standard deviation for over $m=100$ realizations of the predicted solutions of the entire field, and compare them with the noise-free test data. We note how the method is able to accurately detect the spatial dynamics, even when extrapolating on a time window twice longer than the training range.
@@
Figure 4: Results of the offline training a), consisting of the VENI and VINDy steps, for the reaction-diffusion problem. In the VINDy box, the identified posterior distributions of the coefficients are displayed. We note that the mixed linear terms, which are the dominant terms for the observed dynamics, are accurately identified as significant. In the VENI box, we highlight how the method is capable of learning an encoding function from noisy data, resulting in two latent variables that exhibit the expected oscillatory behavior. A comparison of solution fields between the approximation and the numerical reference solution is given in b) for a training $t=15$ and extrapolated time instances $t\in\{20,35\}$, while forward uncertainty quantification at the different reduction levels, performed with the VICI procedure.
3.3 MEMS resonator: Identification of structural mechanical second-order systems
[…]
@@ 3.3 MEMS resonator: Identification of structural mechanical second-order systems
where $\Omega_{0}$ is the domain occupied by the device in the undeformed configuration, described by material coordinates $\varkappa$, and $T=1755\,\mu\text{s}$. Eq. (20a) expresses the conservation of momentum, where $\rho_{0}$ is the initial density, $\mathbf{P}$ is the first Piola-Kirchhoff stress and $\mathbf{B}$ are the external body forces. In this application, we consider a periodic body load proportional to the first vibrational eigenfunction, with the amplitude and frequency given by the parameters $F$ and $\omega$, respectively. The input parameter vector hence becomes $\bm{\beta}~=[\omega,F]^{\intercal}$, defined over a closed and bounded set of dimension two. Eq. (20b) and (20c) define the homogeneous boundary conditions on the Neumann and Dirichlet and boundaries denoted as $\partial\Omega_{N}$ and $\partial\Omega_{D}$, respectively, and $\mathbf{N}$ denotes the outward-directed unit normal on the boundary of $\Omega$. For more details regarding the application, we refer to [13]. A comprehensive summary of network hyperparameters and dataset settings can be found in Table 1-2 in the A.4.
Data generation
We consider $N_{\beta}=56$ parameter instances, corresponding to 28 values for the frequency $\omega$, selected in the range $0.526,0.564]\,\text{rad}/\mu\text{s}$ with a finer sampling around the natural frequency $\omega_{0}$, and for two values of the forcing amplitude $F\in\{0.125,0.250\}\,\mu\text{N}$. In total, $N_{\beta}^{\text{train}}=28$ parameter instances are randomly selected to construct the training set and the remaining $N_{\beta}^{\text{test}}=28$ are retained for testing. Numerical solutions for the displacement are computed for each parameter configuration, by discretizing the system (20) using the finite element method over a spatial mesh consisting of $N_{\text{dof}}=7821$ degrees of freedom as in [13]. Training data are calculated up to $T_{\text{train}}=1091\,{\mu}s<T$ and corrupted by additive noise following a normal distribution whose scale is proportional to the mean displacement level in all training simulations.
@@
As in the previous example, the data (including displacements, velocities, and accelerations) are preliminary projected onto a POD basis of $N=N_{\text{POD}}=3$ modes (extracted from the noisy displacement training data).
@@
The encoder is composed of three hidden layers that use 32 neurons each. The decoder has symmetrical structure. Although the dynamics of this system lies on a two-dimensional manifold in the phase space, we manage to reduce the dimensionality to a single latent variable $n=1$, as the dependence on the velocity is the minimal for the master bending mode that we inspect herein. As in (20), structural mechanics problems are often governed by second-order governing equations. Therefore we model the reduced dynamics with a second-order ODE, by augmenting the reduce states with the first derivatives $\dot{z}$:
$\frac{\text{d}}{\text{dt}}\begin{bmatrix}\;z\;\\ \;\dot{z}\;\end{bmatrix}=\mathbf{f}\left(z,\dot{z};\bm{\beta}\right)\approx\begin{bmatrix}\;\dot{z}\;\\ \;\bm{\Xi}\bm{\Theta}(z,\dot{z};\bm{\beta})\;\end{bmatrix}$. (21)
The corresponding accelerations at latent level $\ddot{\bm{z}}$ can be derived using chain rule as it is done for the velocities $\dot{\bm{z}}$. Note that in case we did not want to exploit any prior knowledge on the second-order structure of the problem, we could have equivalently modeled the problem by considering $n=2$ latent variables and identifying a first-order ODE system as in the previous example. The library $\bm{\Theta}$ features polynomials up to the third degree with respect to both $\bm{z}$ and $\dot{\bm{z}}$. Moreover, since the beam resonator is forced harmonically with amplitude and frequency parameterized by $F$ and $\omega$, we incorporate the parameter dependency by including the library term $F\cos(\omega t)$. We set Gaussian priors on the reduced states $\bm{z},\dot{\bm{z}}]^{\top}\sim\mathcal{N}(\mathbf{0},\bm{I})$ and Laplacian prior on each of the VINDy coefficient distributions $\bm{\Xi}_{ij}\sim\mathcal{L}(0,1)$. Training is performed for 2500 epochs, and the weights of the networks that performed best in terms of total loss on a validation subset of the training data used for inference. After training we apply PDF zero tresholding to only keep the terms that are important to describe the dynamics. In particular, all terms with a $pdf_{\bm{\Xi}_{ij}}(0)>1.9$ are neglected in the model.
Results